OpenStudy (anonymous):

If the price of a stock increases by 40% and then by 25% by what percentage has the price increased from its original value

OpenStudy (anonymous):

let x be the initial price then after 40% it becomes x+0.4x i.e.1.4x then after 25% increase 1.4x+0.25x=1.65x Can u do from now on

OpenStudy (anonymous):

sorry in second step 1.4x+0.25*1.4x

OpenStudy (anonymous):

How did you get 1.4x after you added x+0.4

OpenStudy (anonymous):

I am adding x+0.4x

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

Now 1.4x+0.25*1.4x=1.75x So net increase is 1.75x-x=0.75x 0.75x/x=0.75*100=75%

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

My pleasure as always